# arXiv:2603.00637v1[cond-mat.mtrl-sci] 28 Feb 2026

### v1

@@ Abstract
We report an all-electron implementation of the quasiparticle self-consistent GWGW (QSGWGW) method for molecules and periodic systems within the framework of numerical atomic orbitals (NAOs), as implemented in the LibRPA software package. We present systematic benchmark calculations on molecular systems, as well as a diverse set of periodic systems including typical semiconductors and wide-gap insulators. Our results demonstrate that the present NAO-based QSGWGW workflow yields molecular ionization potentials and quasiparticle band gaps for periodic solids that are consistent with established reference benchmarks, supporting the correctness of the implementation.
@@ I Introduction
In computational chemistry and condensed matter physics, accurately solving the quantum many-body problem from first principles remains a fundamental challenge. While Kohn-Sham (KS) density functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) is the most widely adopted approach, it suffers from well-known limitations of approximate exchange-correlation (XC) functionals, particularly in predicting excited-state properties and describing strongly correlated systems. For a large variety of molecules and materials, many-body perturbation theory within the $GW$ approximation Hedin (1965); Strinati et al. (1982); Hybertsen and Louie (1987); Godby et al. (1988); Aryasetiawan and Gunnarsson (1995); Aulbur et al. (2000); Onida et al. (2002); Rinke et al. (2005); Marom et al. (2012); Golze et al. (2019); Reining (2018); Lange and Berkelbach (2018) offers a satisfactory approach for quantitatively describing single-particle excitations. Within $GW$, the many-body self-energy is given by the product of the Green’s function $G$ and the screened Coulomb interaction $W$, which captures the dominant screening processes in a many-body system and yields physically meaningful quasiparticle energies.
Over the past four decades, the $GW$ method has been developed into a versatile approach encompassing a variety of practical computational schemes, depending on whether and how self-consistency is incorporated. The simplest variant of $GW$ is the so-called $G^{0}W^{0}$ scheme, where the Green’s function $G^{0}$ and screened Coulomb interaction $W^{0}$ are constructed from the outputs of a preceding mean-field calculation. $G^{0}W^{0}$ based on KS-DFT starting points with local or semilocal exchange-correlation functionals, such as the local density approximation (LDA) or generalized gradient approximations (GGAs) Perdew et al. (1996); Onida et al. (2002); Reining (2018), has proven highly successful in predicting quantitatively accurate electronic band gaps for a large variety of materials.
@@
In principle, solving the set of $GW$ equations in a fully self-consistent manner (the so-called scGWGW scheme) is feasible Ku and Eguiluz (2002); Stan et al. (2006); Rostgaard et al. (2010); Caruso et al. (2013, 2012); Kutepov (2016, 2017); Shishkin and Kresse (2007); Grumet et al. (2018a); Holm and von Barth (1998); Koval et al. (2014); Wen et al. (2024); Maggio and Kresse (2017); Kutepov et al. (2009). In the fully self-consistent limit, scGWGW is conserving in the sense of Baym and Kadanoff and removes the explicit starting-point dependence Baym and Kadanoff (1961); Baym (1962). However, scGWGW is computationally demanding and, without vertex corrections, often overestimates band gaps for semiconductors and insulators Shishkin and Kresse (2007); Shishkin et al. (2007). Moreover, neither $G^{0}W^{0}$ nor scGWGW provides an updated single-particle Hamiltonian with explicit quasiparticle wavefunctions, which restricts their utility in theoretical analyses requiring orbital information, such as investigations of topological materials. These limitations motivate the development of an alternative approach that can deliver quasiparticle solutions while maintaining theoretical rigor. The quasiparticle self-consistent GWGW (QSGWGW) method, developed by Faleev, van Schilfgaarde, and Kotani Faleev et al. (2004); van Schilfgaarde et al. (2006); Kotani et al. (2007), fulfills these requirements.
@@
The fundamental principle of QSGWGW is to approximate the dynamic self-energy $\Sigma(\omega)$ by a static, Hermitian XC potential $V^{\mathrm{xc}}$. This potential is used to iteratively update a static mean-field Hamiltonian until self-consistency is reached, enabling the simultaneous optimization of quasiparticle energies and wavefunctions. Different prescriptions exist for constructing the static, Hermitian $V^{\mathrm{xc}}$ from $\Sigma(\omega)$, including the commonly used “Mode A”/“Mode B” constructions van Schilfgaarde et al. (2006) and SRG-based approaches Marie and Loos (2023). Moreover, QSGWGW can be systematically improved by incorporating vertex corrections (e.g., ladder diagrams in $W$) Cunningham et al. (2023); Kutepov (2022); Förster (2025); Sakakibara et al. (2020); Förster and Visscher (2022a) and lattice-polarization effects on $W$ Lambrecht et al. (2017). In this work, we adopt the standard formulation proposed by the original authors Faleev et al. (2004); van Schilfgaarde et al. (2006); Kotani et al. (2007). The reliability and broad applicability of this approach have been extensively validated across a diverse range of materials, from conventional semiconductors to complex transition metal oxides Das et al. (2015); Salas-Illanes et al. (2022).
@@
Naturally, the performance of the QSGWGW method is closely tied to the underlying numerical implementation. Over the years, QSGWGW has been implemented within various numerical frameworks, including linear muffin-tin orbitals (LMTO) Kotani et al. (2007), augmented-plane-wave based all-electron methods Deguchi et al. (2016); Salas-Illanes et al. (2022), Gaussian-type orbitals (GTOs)Lei and Zhu (2022), and more recently on graphics processing units (GPUs) Obata et al. (2025). Closely related self-consistent $GW$ implementations in plane-wave projector augmented-wave (PAW) frameworks are also well established Shishkin and Kresse (2006, 2007).
@@
In recent years, numerical atomic orbitals (NAOs) have emerged as a powerful basis-set framework for electronic structure calculations. NAOs offer several key advantages. First, the strict localization of NAO basis functions naturally facilitates the description of localized states—such as core electrons and strongly correlated $d/f$ electrons—enabling all-electron calculations without the need for pseudopotentials. Second, the compactness of NAOs significantly reduces the dimensionality of the Hilbert space. Unlike plane-wave methods, which require thousands of basis functions and a large number of unoccupied states to converge the screened interaction, NAOs typically require only tens to hundreds of functions, thereby greatly alleviating the slow convergence with respect to the number of unoccupied states. Furthermore, when combined with the resolution of identity (RI) technique Feyereisen et al. (1993); Vahtras et al. (1993); Weigend et al. (1998); Förster and Visscher (2020), two-electron integrals within the NAO framework can be efficiently decomposed, drastically reducing memory requirements and improving computational scaling. Benchmarks using established NAO codes, such as FHI-aims, demonstrate that the accuracy achieved on the GW100 test set van Setten et al. (2015) is comparable to that of plane-wave implementations, while the compactness of NAO basis sets and low-scaling formulations can offer favorable efficiency Ren et al. (2012); Förster and Visscher (2020); Ren et al. (2021a).
@@
Despite the existence of molecular and periodic $G^{0}W^{0}$ implementations based on NAOs Ren et al. (2012, 2021a); Zhang et al. (2026), an all-electron QSGWGW implementation within the NAO framework has not yet been reported. This gap has hindered the application of NAO-based methods to complex systems where self-consistency is essential, such as topological materials and strongly correlated oxides.
Developing such an implementation is therefore highly desirable, as it would combine the computational efficiency of NAOs with the rigorous quasiparticle framework of QSGWGW.
@@
In this paper, we present an implementation of the QSGWGW method based on all-electron NAOs within the LibRPA software package, applicable to both molecules and periodic systems. We systematically benchmark the performance of our implementation for the vertical ionization potentials (IPs) of small molecules from the GW100 test set, as well as the band gaps of a diverse range of semiconductors and insulators. Furthermore, we present a detailed comparison with the $G^{0}W^{0}$ method and other QSGWGW implementations to comprehensively evaluate the accuracy and reliability of our approach.
@@
This paper is organized as follows. In Sec. II, we outline the theoretical framework, detailing the QSGWGW formalism and its implementation using the NAO-based real-space formalism combined with the localized resolution of identity (LRI) technique. Sec. III discusses specific implementation details in LibRPA, including the self-consistency workflow, convergence behavior, and an analysis of the basis dependence in the analytic continuation process. Sec. IV presents benchmark results for molecular systems (Sec. IV.1) and periodic solids (Sec. IV.2), comparing our results with standard DFT, $G^{0}W^{0}$, and other QSGWGW implementations. Finally, we summarize our findings and provide an outlook in Sec. V.
@@ II Theoretical Framework
This section outlines the theoretical foundations of the QSGWGW method. We begin with a review of the $GW$ approximation and the one-shot $G^{0}W^{0}$ approach, followed by a detailed formulation of the QSGWGW scheme. We then present the NAO-based formalism combined with the localized resolution of identity (LRI) technique, which forms the algorithmic basis of our implementation.
@@ II.1 The $GW$ approximation and the $G^{0}W^{0}$ scheme
The interacting Green’s function satisfies the Dyson equation, which relates it to the non-interacting Green’s function $G^{0}$ via the self-energy $\Sigma$:
@@
where $W^{0}$ is the screened Coulomb interaction calculated typically within the random phase approximation (RPA).
@@
where $\varepsilon^{-1}$ is the inverse of the dielectric function $\varepsilon$, which describes the screening process and can be computed as
@@
Here, $v(\mathbf{r},\mathbf{r}^{\prime})=1/|\mathbf{r}-\mathbf{r}^{\prime}|$ is the bare Coulomb interaction and $\chi^{0}$ is the irreducible polarizability. Under the RPA, $\chi^{0}$ equals the non-interacting density response function that is given by
@@
As is evident from the above equations, in the $G^{0}W^{0}$ approach, the calculation relies entirely on the non-interacting Green’s function $G^{0}$, which is typically constructed from a preceding DFT calculation:
$G^{0}(\mathbf{r},\mathbf{r}^{\prime},\omega)=$ $\sum_{p}(1-f_{p})\frac{\psi_{p}(\mathbf{r})\psi_{p}^{*}(\mathbf{r}^{\prime})}{\omega+\epsilon_{F}-\epsilon_{p}+\mathrm{i}\eta}$ (6)
$+\sum_{p}f_{p}\frac{\psi_{p}(\mathbf{r})\psi_{p}^{*}(\mathbf{r}^{\prime})}{\omega+\epsilon_{F}-\epsilon_{p}-\mathrm{i}\eta}$
where $\psi_{p}(\mathbf{r})$ and $\epsilon_{p}$ are the KS orbitals and orbital energies, respectively, $f_{p}$ is the occupation number, $\epsilon_{F}$ is the Fermi energy (chemical potential), and $\eta$ is a positive infinitesimal.
Within the $G^{0}W^{0}$ approach, the quasiparticle excitation energies are obtained by solving the quasiparticle equation
@@
where the self-energy depends on the quasiparticle energy itself. This nonlinear equation is solved self-consistently for each state, typically by root-finding algorithms. Importantly, in $G^{0}W^{0}$, the Green’s function $G^{0}$ and screened interaction $W^{0}$ are fixed from the initial DFT calculation and are not updated during this process—hence the term “one-shot” $GW$. Only the diagonal elements of the self-energy matrix are required.
@@ II.2 The QSGWGW scheme
$+\sum_{p\neq p^{\prime}}\mathinner{|{\psi_{p}}\rangle}\mathrm{Re}[\Sigma^{GW}(\epsilon_{F})]_{pp^{\prime}}\mathinner{\langle{\psi_{p^{\prime}}}|}.$
Here $\epsilon_{p}$ is the $p$-th eigenvalue of the Hamiltonian $H^{\mathrm{QS}GW}$, $\epsilon_{F}$ is the Fermi energy, and $\mathrm{Re}$” denotes the Hermitian part of the self-energy matrix, i.e., $\mathrm{Re}[A]_{pp^{\prime}}\equiv(A_{pp^{\prime}}+A_{p^{\prime}p}^{*})/2$. Formally, the construction in “Mode A” can also be derived by minimizing the length of the gradient of the Klein total energy functional over non-interacting Green’s functions (given by $(\omega-H^{\mathrm{QS}GW})^{-1}$), as demonstrated by Ismail-Beigi Ismail-Beigi (2017).
In a practical self-consistent QSGWGW calculation, the following steps need to be executed iteratively:
@@
The above-outlined QSGWGW scheme has been implemented in the LibRPA package Shi et al. (2025); Zhang et al. (2026). LibRPA was originally developed for the RPA correlation energy Shi et al. (2025) and one-shot $G^{0}W^{0}$ calculations Zhang et al. (2026). LibRPA is interfaced with FHI-aims to acquire necessary inputs from DFT calculations. We use the NAO-based real-space formalism combined with the localized resolution of identity (LRI) technique to construct the XC potential operator of QSGWGW efficiently, which will be detailed in the next subsection.
II.3 NAO-based formalism for QSGWGW
To implement the QSGWGW formalism efficiently within the NAO framework, we employ a space-time approach Ren et al. (2021a); Zhang et al. (2026) combined with the LRI technique Lin et al. (2020). This approach facilitates the computation of the full $GW$ self-energy matrix required by the QSGWGW scheme. Furthermore, the use of NAO basis functions and the LRI technique gives rise to significant sparsity in the LRI expansion coefficients. Exploiting this sparsity leads to an $\mathcal{O}(N^{2})$-scaling algorithm for constructing the $GW$ self-energy, compared to the $\mathcal{O}(N^{4})$ scaling in the conventional approach. A detailed description of the underlying $G^{0}W^{0}$ implementation has been provided in a separate work Zhang et al. (2026). Here, we only summarize the key steps and equations specific to the QSGWGW implementation.

### v2

# All-electron Quasiparticle Self-consistent $GW$ for Molecules and Periodic Systems within the Numerical Atomic Orbital Framework

@@ Abstract
We report an all-electron implementation of the quasiparticle self-consistent GWGW (QSGWGW) method for molecular and periodic systems within the framework of numerical atomic orbitals (NAOs), as implemented in the LibRPA software package. Our implementation is based on the space-time formalism, combined with the localized resolution-of-identity approximation to treat two-electron quantities. We found that analytical continuation of the self-energy matrix, in combination with the “Mode B” QSGWGW scheme, can yield stable self-consistent quasiparticle energy spectra. Systematic benchmark calculations on molecules and crystalline solids (including typical semiconductors and wide-gap insulators) demonstrate that our NAO-based QSGWGW scheme yields molecular ionization potentials and quasiparticle band gaps for periodic solids that are consistent with reference results from established implementations. Our work opens the way for large-scale QSGWGW calculations, taking advantage of the NAO-based low-scaling algorithm previously developed for the $G^{0}W^{0}$ method.
@@ I Introduction
In computational chemistry and condensed matter physics, accurately solving the quantum many-body problem from first principles remains a fundamental challenge. While Kohn-Sham (KS) density functional theory (DFT) Hohenberg and Kohn (1964); Kohn and Sham (1965) is the most widely adopted approach, it suffers from well-known limitations of approximate exchange-correlation (XC) functionals, particularly in predicting excited-state properties and describing strongly correlated systems. For a large variety of molecules and materials, many-body perturbation theory within the $GW$ approximation Hedin (1965); Strinati et al. (1982); Hybertsen and Louie (1987); Godby et al. (1988); Aryasetiawan and Gunnarsson (1995); Aulbur et al. (2000); Onida et al. (2002); Rinke et al. (2005); Marom et al. (2012); Golze et al. (2019); Reining (2018); Lange and Berkelbach (2018) offers a powerful approach for quantitatively describing single-particle excitations. Within $GW$, the many-body self-energy is given by the product of the Green’s function $G$ and the screened Coulomb interaction $W$, which captures the dominant screening effect in a many-body system and yields physically meaningful quasiparticle energies.
Over the past four decades, the $GW$ method has been developed into a versatile approach encompassing a variety of practical computational schemes, depending on whether and how self-consistency is incorporated. The simplest variant of $GW$ is the so-called $G^{0}W^{0}$ scheme, where the Green’s function $G^{0}$ and screened Coulomb interaction $W^{0}$ are constructed from the outputs of a preceding mean-field calculation. $G^{0}W^{0}$ based on KS-DFT starting points with local or semilocal exchange-correlation functionals, such as the local density approximation (LDA) or generalized gradient approximations (GGAs) Perdew et al. (1996); Hybertsen and Louie (1987); Onida et al. (2002); Reining (2018), has proven highly successful in predicting quantitatively accurate electronic band gaps for a large variety of materials.
@@
In principle, solving the set of $GW$ equations in a fully self-consistent manner (the so-called scGWGW scheme) is feasible Ku and Eguiluz (2002); Stan et al. (2006); Rostgaard et al. (2010); Caruso et al. (2013, 2012); Kutepov (2016, 2017); Shishkin and Kresse (2007); Grumet et al. (2018); Holm and von Barth (1998); Koval et al. (2014); Wen et al. (2024); Maggio and Kresse (2017); Kutepov et al. (2009). In the fully self-consistent limit, scGWGW is conserving in the sense of Baym and Kadanoff and removes the explicit starting-point dependence Baym and Kadanoff (1961); Baym (1962). However, scGWGW is computationally demanding and, without vertex corrections, often overestimates band gaps for semiconductors and insulators Shishkin and Kresse (2007); Shishkin et al. (2007). Moreover, neither $G^{0}W^{0}$ nor scGWGW provides an updated single-particle Hamiltonian with explicit quasiparticle wavefunctions, which limits their utility in theoretical analyses that require orbital information, such as the investigation of topological materials. These limitations motivate the development of an alternative approach that can deliver quasiparticle solutions while maintaining theoretical rigor. The quasiparticle self-consistent GWGW (QSGWGW) method, developed by Faleev, van Schilfgaarde, and Kotani Faleev et al. (2004); van Schilfgaarde et al. (2006); Kotani et al. (2007), fulfills these requirements.
@@
The fundamental principle of QSGWGW is to approximate the dynamic self-energy $\Sigma(\omega)$ by a static, Hermitian XC potential $V^{\mathrm{xc}}$. This potential is used to iteratively update a static mean-field Hamiltonian until self-consistency is reached, enabling the simultaneous optimization of quasiparticle energies and wavefunctions. Different prescriptions exist for constructing the static, Hermitian $V^{\mathrm{xc}}$ from $\Sigma(\omega)$, including the commonly used “Mode A”/“Mode B” constructions van Schilfgaarde et al. (2006) and similarity renormalization group (SRG)-based approaches Marie and Loos (2023). Moreover, QSGWGW can be systematically improved by incorporating vertex corrections (e.g., ladder diagrams in $W$) Cunningham et al. (2023); Kutepov (2022); Förster (2025); Sakakibara et al. (2020); Förster and Visscher (2022a) and lattice-polarization effects on $W$ Lambrecht et al. (2017). In this work, we adopt the standard formulation proposed by the original authors Faleev et al. (2004); van Schilfgaarde et al. (2006); Kotani et al. (2007). The reliability and broad applicability of this approach have been extensively validated across a diverse range of materials, from conventional semiconductors to complex transition metal oxides Das et al. (2015); Salas-Illanes et al. (2022).
@@
Naturally, the computational efficiency of the QSGWGW method is closely tied to the underlying numerical implementation. Over the years, QSGWGW has been implemented within various numerical frameworks, including linear muffin-tin orbitals (LMTO) Kotani et al. (2007), augmented-plane-wave based all-electron methods Deguchi et al. (2016); Salas-Illanes et al. (2022), and Gaussian-type orbitals (GTOs) Lei and Zhu (2022). More recently, implementations on GPU architectures have also been reported to accelerate the QSGWGW workflow Obata et al. (2025).
@@
In recent years, numerical atomic orbitals (NAOs) have emerged as a powerful basis-set framework for electronic structure calculations. NAOs offer several key advantages: First, the flexible shape of NAO basis functions naturally facilitates the description of localized states—such as core electrons and strongly correlated $d/f$ electrons—enabling all-electron calculations that are free of pseudopotential approximations. Second, the compactness of NAOs significantly reduces the dimensionality of the Hilbert space. Unlike plane-wave methods, which require thousands of basis functions and a large number of unoccupied states to converge the screened Coulomb interaction and the self-energy, NAOs typically require only tens to hundreds of functions, thereby greatly alleviating the slow convergence with respect to the number of unoccupied states. Furthermore, when combined with the resolution of identity (RI) technique Feyereisen et al. (1993); Vahtras et al. (1993); Weigend et al. (1998); Förster and Visscher (2020), two-electron integrals within the NAO framework can be efficiently decomposed, drastically reducing memory requirements and improving computational scaling. Benchmarks using established NAO codes, such as FHI-aims, demonstrate that the accuracy achieved on the GW100 test set van Setten et al. (2015) is comparable to that of plane-wave implementations, while the compactness of NAO basis sets and low-scaling formulations can offer favorable efficiency Ren et al. (2012); Förster and Visscher (2020); Ren et al. (2021); Shi et al. (2024).
@@
Despite the existence of molecular and periodic $G^{0}W^{0}$ implementations based on NAOs Ren et al. (2012, 2021); Zhang et al. (2026), an all-electron QSGWGW implementation within the NAO framework has not yet been reported. The absence of the QSGWGW feature has hindered the application of NAO-based methods to complex systems where self-consistency is essential, such as topological materials and strongly correlated oxides. Developing such an implementation is therefore highly desirable, as it would combine the computational efficiency of NAOs with the rigorous quasiparticle framework of QSGWGW.
@@
In this paper, we present an implementation of the QSGWGW method based on all-electron NAOs within the LibRPA software package, applicable to both molecules and periodic systems. We systematically benchmark the performance of our implementation for the vertical ionization potentials (IPs) of small molecules from the GW100 test set, as well as the band gaps of a diverse set of semiconductors and insulators. Furthermore, we present a detailed comparison with the $G^{0}W^{0}$ method and other QSGWGW implementations to comprehensively evaluate the accuracy and reliability of our approach.
@@
This paper is organized as follows. In Sec. II, we outline the theoretical framework, detailing the QSGWGW formalism and its implementation using the NAO-based real-space formalism combined with the localized resolution of identity (LRI) technique. Sec. III discusses specific implementation details in LibRPA, including the self-consistency workflow, the convergence behavior, and an analysis of the basis dependence in the analytic continuation process. Sec. IV presents benchmark results for molecular systems (Sec. IV.1) and periodic solids (Sec. IV.2), comparing our results with those obtained with standard DFT, $G^{0}W^{0}$, and other QSGWGW implementations. Finally, we summarize our findings and provide an outlook in Sec. V.
@@ II Theoretical Framework
This section outlines the theoretical foundations of the QSGWGW method. We begin with a review of the $GW$ approximation and the one-shot $G^{0}W^{0}$ approach, followed by a detailed formulation of the QSGWGW scheme. We then present the NAO-based formalism combined with the LRI technique, which forms the algorithmic basis of our implementation.
@@ II.1 The $GW$ approximation and the $G^{0}W^{0}$ scheme
The interacting Green’s function $G$ satisfies the Dyson equation, which relates $G$ to the non-interacting Green’s function $G^{0}$ via the self-energy $\Sigma$:
@@
where $W^{0}$ is the screened Coulomb interaction, typically evaluated within the random phase approximation (RPA).
@@
where $\varepsilon^{-1}$ is the inverse of the dielectric function $\varepsilon$, which describes the electron screening process and can be computed as
@@
Here, $v(\mathbf{r},\mathbf{r}^{\prime})=1/|\mathbf{r}-\mathbf{r}^{\prime}|$ is the bare Coulomb interaction and $\chi^{0}$ is the irreducible polarizability. Under the RPA, $\chi^{0}$ equals the non-interacting density response function, given by
@@
As is evident from the above equations, in the $G^{0}W^{0}$ approach, the calculation relies entirely on the non-interacting Green’s function $G^{0}$, which is usually constructed from a preceding DFT calculation:
$G^{0}(\mathbf{r},\mathbf{r}^{\prime},\omega)=$ $\sum_{p}(1-f_{p})\frac{\psi_{p}(\mathbf{r})\psi_{p}^{*}(\mathbf{r}^{\prime})}{\omega+\epsilon_{\rm F}-\epsilon_{p}+\mathrm{i}\eta}$ (6)
$+\sum_{p}f_{p}\frac{\psi_{p}(\mathbf{r})\psi_{p}^{*}(\mathbf{r}^{\prime})}{\omega+\epsilon_{\rm F}-\epsilon_{p}-\mathrm{i}\eta}$
where $\psi_{p}(\mathbf{r})$ and $\epsilon_{p}$ are the KS orbitals and orbital energies, respectively, $f_{p}$ is the occupation number, $\epsilon_{\rm F}$ is the Fermi energy (chemical potential), and $\eta$ is a positive infinitesimal.
Within the $G^{0}W^{0}$ scheme, the quasiparticle excitation energies are obtained by solving the quasiparticle equation
@@
where the self-energy depends on the quasiparticle energy itself. This nonlinear equation is solved self-consistently for each state, typically by root-finding algorithms. Importantly, in $G^{0}W^{0}$, the Green’s function $G^{0}$ and screened interaction $W^{0}$ are fixed from the initial DFT calculation and are not updated during the calculation — hence the term “one-shot” $GW$. Only the diagonal elements of the self-energy matrix are required.
@@ II.2 The QSGWGW scheme
$+\sum_{p,p\neq p^{\prime}}\mathinner{|{\psi_{p}}\rangle}\mathrm{Re}[\Sigma^{GW}(\epsilon_{\rm F})]_{pp^{\prime}}\mathinner{\langle{\psi_{p^{\prime}}}|}.$
Here $\epsilon_{p}$ is the $p$-th eigenvalue of the Hamiltonian $H^{\mathrm{QS}GW}$, $\epsilon_{\rm F}$ is the Fermi energy, and $\mathrm{Re}$” denotes the Hermitian part of the self-energy matrix, i.e., $\mathrm{Re}[A]_{pp^{\prime}}\equiv(A_{pp^{\prime}}+A_{p^{\prime}p}^{*})/2$. Formally, the construction in “Mode A” can also be derived by minimizing the length of the gradient of the Klein total energy functional over non-interacting Green’s functions (given by $(\omega-H^{\mathrm{QS}GW})^{-1}$), as demonstrated by Ismail-Beigi Ismail-Beigi (2017).
In a practical QSGWGW calculation, the following steps need to be executed:
@@
The above-outlined QSGWGW scheme has been implemented in the LibRPA package Shi et al. (2025); Zhang et al. (2026). LibRPA was originally developed for the RPA correlation energy Shi et al. (2025) and one-shot $G^{0}W^{0}$ calculations Zhang et al. (2026). LibRPA is interfaced with FHI-aims to acquire necessary inputs from DFT calculations. We use the NAO-based real-space formalism combined with the LRI technique to construct the XC potential operator of QSGWGW efficiently, which will be detailed in the next subsection.
II.3 NAO-based formalism for QSGW
To implement the QSGWGW formalism efficiently within the NAO framework, we employ a space-time approach Ren et al. (2021); Zhang et al. (2026) combined with the LRI technique Lin et al. (2020). This approach facilitates the computation of the full $GW$ self-energy matrix required by the QSGWGW scheme. Furthermore, the use of NAO basis functions and the LRI technique gives rise to significant sparsity in the LRI expansion coefficients. Exploiting this sparsity leads to an $\mathcal{O}(N^{2})$-scaling algorithm for constructing the $GW$ self-energy, compared to the $\mathcal{O}(N^{4})$ scaling in the conventional approach. A detailed description of the underlying $G^{0}W^{0}$ implementation has been provided in a separate work Zhang et al. (2026). Here, we only summarize the key steps and equations specific to the QSGWGW implementation.